OpenStudy (anonymous):

find the area between the curves f(x)=e^0.8x and g(x)=-2^0.5x +1 for the interval [-3,2]

OpenStudy (amistre64):

f(x)-g(x) sounds familiar to me

OpenStudy (amistre64):

in other words, the height of each partition is simply the distance from f(x) to g(x)

OpenStudy (anonymous):

can you use a calculator or all by hand?

OpenStudy (amistre64):



OpenStudy (anonymous):

i need to do it by hand

OpenStudy (anonymous):

you will need to know which function is above and which is below.

OpenStudy (anonymous):

Integrate top function minus the bottom function.

OpenStudy (anonymous):

on the given interval.

OpenStudy (amistre64):

\[\int f(x)-g(x)dx\] no you wont, all that changes is a sign and area is always positive for drop a sign

OpenStudy (amistre64):

\[\int_{-3}^{2} e^{0.8x}-(-2^{0.5x} +1)dx\] hmm, might need to write that out better

OpenStudy (amistre64):

or is that right?

OpenStudy (amistre64):

you might wanna check to see that nothing cross tho

OpenStudy (amistre64):

if it does then you gotta split it up at the intersection

OpenStudy (amistre64):

does f(x)=g(x) at any point in the interval?

OpenStudy (anonymous):

well i dont think so

OpenStudy (amistre64):

if i got your equations right, the wolf says they cross

OpenStudy (anonymous):

i figured out the integral thing but I didn't get the answer 6.935

OpenStudy (anonymous):

Let me say again that the graph is important. On this interval the graphs intersect and therefore the integral must be seperated.

OpenStudy (amistre64):

yes, intersections are important; height? not so much :)

OpenStudy (anonymous):

(-1.282, 0.359) is the intersection right?

OpenStudy (anonymous):

You will integrate from -3 to the point of intersection with g(x) - f(x) and then from the point of intersection to 2 with f(x)-g(x)....

OpenStudy (amistre64):

\[\int_{-3}^{i} e^{0.8x}+2^{0.5x} -1\ dx+\int_{i}^{2} e^{0.8x}+2^{0.5x} -1\ dx\] where i is the intersection of f and g

OpenStudy (anonymous):

I thought you mentioned solving by hand?

OpenStudy (amistre64):

and if one of those is negative, then toss out the "-" sign :)

OpenStudy (anonymous):

yes I need to solve it by hand, and I did what you just said but i got the wrong answer

OpenStudy (anonymous):

....can you show your work?

OpenStudy (amistre64):

the wolf agrees that -1.282 is the intersect

OpenStudy (anonymous):

That is correct....but that is not solving by hand as I stated earlier.

OpenStudy (anonymous):

\[\int\limits_{-3}^{-1.282} -2^{0.5x}-e ^{0.8x} dx + \int\limits_{-1.282}^{2} e ^{0.8x}-(-2^{0.5x}) dx\]

OpenStudy (amistre64):

its checking if her results are accurate tho.

OpenStudy (amistre64):

wheres the "1"?

OpenStudy (amistre64):

you might be better of with an exact intersect ....

OpenStudy (amistre64):

and who is making you do this by hand?

OpenStudy (amistre64):

its one thing to understand a concept; its another to torture ....

OpenStudy (anonymous):

No...It is important to understand the concept and to use technology that is required in the classroom and on ap exams... I can help her learn to enter the information into the calculator and how she must set it up by hand to write the integral that represents the area. Dropping a sign because area is not negative is not good enough reasoning for a solution.

OpenStudy (anonymous):

\[[-2*2^{0.5x}- 5/4e ^{0.8x}] + [5/4 e ^{0.8x}+2*2^{0.5x}]\]

OpenStudy (anonymous):

Have a good night...

OpenStudy (amistre64):

oh but it is; since 5-3 = 2, and 3-5 = -2 all that changes is sign, not abs value

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

youre still missing that +1 from the top that ints up into an x

OpenStudy (anonymous):

oh yess. I see may be that's why i dont get the answer

OpenStudy (amistre64):

most likely :) the rest of it seems like youve got a pretty good handle on it

OpenStudy (anonymous):

how can i forget that "1" OMG i'm so stupid

OpenStudy (amistre64):

it happens :)

OpenStudy (anonymous):

ok i'll try again. thank you so muchhhhhhhh and have a great night :)